OpenStudy (anonymous):

Which two points satisfy y = -x2 + 2x + 4 and x + y = 4? (0, 4) and (4, 0) (0, 4) and (3, 1) (7, -3) and (0, 4) (7, -3) and (4, 0)

OpenStudy (mathmale):

Hello, zach! First of all, please express your exponentiation using the " ^ " symbol: Correct: y = -x^2 + 2x + 4 Incorrect: y = -x2 + 2x + 4

OpenStudy (mathmale):

Among the several methods you have to choose from in solving this problem: 1. Substitute the coordinates of each given point into both of the given equations. If both of y our equations come out to be true, you've found the solution. 2. solve the system of two equations yourself. Use the substitution method. Compare your results to the four given answer choices. 3. Graph the 2 given equations.

OpenStudy (mathmale):

Note: Choose just one method to begin with.